OpenStudy (anonymous):

18-6k+6(1+3k) solve for k show all work please help mee !

OpenStudy (anonymous):

You would start by distributing first. Distribute 6(1+3k)

OpenStudy (anonymous):

What does everything equal to?

OpenStudy (anonymous):

12(k+2)

OpenStudy (anonymous):

Since it is solving for k, what is the equation equal to? 18-6k+(1+3k)=?

OpenStudy (anonymous):

I have no idea

OpenStudy (anonymous):

Is that the whole problem?

OpenStudy (anonymous):

You can simplify everything, but it will not solve for k

OpenStudy (anonymous):

um wait no its -18-6k=6(1+3k)

OpenStudy (anonymous):

Ok, that makes sense.

OpenStudy (anonymous):

Distribute 6(1+3k)

OpenStudy (anonymous):

You should end up with -18-6k=6+12k

OpenStudy (anonymous):

Subtract 6 from both sides. -24-6k=12k. Then add 6k to both sides. -24=18k. Divide by 18 both sides. -1.3=k or k=-1.3

OpenStudy (anonymous):

Ok thanks !